Find the slope and y- intercept of the line represented by the following equations.



problem 12:


2x- 3y = 6

Answer by fractalier(6550)

From 2x - 3y = 6 we put it in y = mx + b form and get
-3y = -2x + 6
y = (2/3)x - 2
so our slope is 2/3 and the y-intercept is -2